the company applies overhead to jobs at a rate of $40 per direct labor hour. during the year the company used a total of 11,628 direct labor hours. the company budgeted overhead of 506,415. Actual overhead was 539,737. By how much was overhead underapplied?
Answer #1

Applied overheads = Overhead rate *Direct labor Hours = 40*11,628 = 465,120

Overheads Underapplied = 539,737-465,120 = $74,617 Under applied (Answer)